# FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding

### v2

@@ 4 Methodology and Experimental Setup
To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (Xia et al., 2024) with continual learning (Kirkpatrick et al., 2017). This section offers a thorough overview of the training procedure and evaluation setup.
@@ 4.1 Training Procedure
Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset. However, as highlighted in studies by (Wang et al., 2023; Li et al., 2024), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance. To mitigate this challenge, Chain of LoRA (CoLA) (Xia et al., 2024), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method). This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module. Xia et al. (2024) shows that this process allows the model to learn higher-rank adaptations more effectively. Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting. Kalajdzievski (2024) observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
In this study we propose Continual-Chain of LoRA (Co-CoLA), an extension of CoLA framework which incorporates rehearsal with replay during training. More specifically, rehearsal training is an approach within the continual learning framework that involves revisiting a portion of previously learned tasks during training new tasks. Despite the limited success of continual learning frameworks, the study by Scialom et al. (2022) demonstrated that continual training of language models, such as T0 Sanh et al. (2022) with rehearsal, can effectively help them in comprehending new instruction via instruction composition, resulting in better generalization and improved performance on new tasks.
@@
The core mathematical operation in LoRA involves updating the low-rank matrices $A$ and $B$, which are applied to modify the transformer layers of the model. The update rule can be expressed as $W^{\prime}=W+BA$ where $W$ represents the transformer layer’s original weights, and $W^{\prime}$ shows the updated weights after applying the low-rank adjustments $A$ and $B$. Essentially, Co-CoLA structures this training procedure into an iterative three phases:
@@
Tuning: In this phase, following the standard LoRA, the base model weights remain frozen, while only the model’s LoRA parameters (represented by matrices $A$ and $B$) are fine-tuned. Additionally, a subset of previously trained data is replayed along with the new data. Formally, given the sequence $T=(T_{1},\ldots,T_{n})$ where $T_{i}$ represents the training data after applying an individual template, the training data augmented with rehearsal is defined as:
$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$ (1)
@@
where r is the rehearsal hyper-parameter that controls the percentage of examples sampled from previous templates $T_{1},\ldots,T_{n}$.
@@
Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights. These merged weights are fixed and do not receive any gradient update in subsequent steps.
@@
Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$). Following Hu et al. (2021) $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
@@ 4.2 Evaluation Setup
Evaluation Tasks: Our model’s performance was evaluated through two categories of task types: those that were part of the training dataset ("Held in") and those introduced to the model for the first time during evaluation ("Held out"). The evaluation dataset encompasses three distinct types of tasks: Sentiment analysis and Query paraphrasing, classified as “Held in” tasks and Textual Entailment which is categorized as a “Held out” task. As illustrated in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, alongside two datasets dedicated to entailment tasks.
Evaluation Metric: To assess the performance of our model relative to several baseline models, we utilized the ROUGE-L metric, which measures the overlap of n-grams between the generated text and reference texts. Specifically, we concentrated on the F1-scores of ROUGE-L, a metric that integrates precision and recall to provide a balanced evaluation. As demonstrated by Wang et al. (2022), the rankings produced by this metric exhibit a strong correlation with accuracy for categorization templates.
@@ 5 Results
To investigate the applicability of FarsInstruct, we choose the Ava model and instruction-tune it using the Co-CoLA framework on a diverse set of templates. Our results were compared against a series of monolingual and multilingual instruction-tuned models and to effectively assess the performance of our model we conduct both quantitative and linguistic evaluations. For a comprehensive overview of the training configuration, please refer to the Appendix A.
Task Type Model ROUGE-L
parsinlu query paraphrasing Held In Aya-13B 45.58
PersianMind-7B 17.07
Mistral-7B 6.89
Dorna-8B 3.85
Ava-8B 6.67
Ava-LoRA-8B 8.73
Co-CoLA-8B 45.86
Digikala Sentiment Analysis Held In Aya-13B 28.41
PersianMind-7B 18.19
Mistral-7B 2.46
Dorna-8B 2.42
Ava-8B 8.69
Ava-LoRA-8B 5.72
Co-CoLA-8B 40.87
FarsTail Held Out Aya-13B 37.61
PersianMind-7B 17.05
Mistral-7B 5.74
Dorna-8B 4.81
Ava-8B 12.48
Ava-LoRA-8B 9.07
Co-CoLA-8B 36.35
Parsinlu Entailment Held Out Aya-13B 42.64
PersianMind-7B 4.45
Mistral-7B 4.93
Dorna-8B 3.32
Ava-8B 15.04
Ava-LoRA-8B 7.18
Co-CoLA-8B 55.32
Table 2: ROUGE-L F1 Scores for Different Models across Tasks
5.1 Quantitative Evaluation
We evaluate our model against several existing models fine-tuned on instruction-specific data. Specifically, PersianMind (University of Tehran, 2024) is a Llama-2 7B based model, trained in 3 phases on different Persian datasets. Though their training data is unavailable, Dorna (PartAI, 2024) and Ava (Moghadam, 2024) are newly introduced models, fine-tuned on the Llama-3 8B model for Persian tasks. Aya (CohereForAI, 2024) is a 13B encoder-decoder model trained on a subset of 25 million samples from the Aya dataset and Mistral-7B (MistralAI, 2024) is a decoder-only model trained on publicly available prompted datasets
Table 2 summarizes the comparative performance of various models, including our proposed method, Co-CoLA, across several NLP Datasets: ParsiNLU Query Paraphrasing, Digikala Sentiment Analysis, FarsTail, and ParsiNLU Entailment. These models are evaluated using ROUGE-L F1 scores. As illustrated in Table 2, Co-CoLA performs comparably well to the Aya model, despite having fewer parameters and being trained on less instruction data and significantly outperforms all other models, indicating the effectiveness of Co-CoLA. The factors contributing to this performance gap are further discussed in Section 6. Moreover, the scores of Ava-LoRA, reflecting the performance of raw LoRA fine-tuning of Ava on FarsInstruct, are inferior to those achieved with Co-CoLA training, highlighting the effectiveness of our method.
Figure 5: Comparative performance of different models on Persian language tasks using the ROUGE-L metric. The bar chart depicts the superior performance of Co-CoLA across multiple tasks, particularly excelling in the ParsiNLU Entailment task.
5.2 Linguistic Evaluation
Our comprehensive linguistic evaluation aimed to further substantiate the effectiveness of Co-CoLA in handling the nuances of the Persian language, compared to the baseline model Ava. The evaluation specifically focused on analyzing the models’ capabilities in terms of coherence, relevance, and linguistic quality, which are critical for assessing the applicability of language models in real-world scenarios.
5.2.1 Evaluation Setup
The evaluation involved detailed analysis by language experts who assessed the output from both models based on predefined criteria. This approach ensures an unbiased evaluation of the models’ performance in generating contextually appropriate and linguistically accurate content.
5.2.2 Evaluation Criteria
The linguistic outputs were evaluated based on three main criteria:
Coherence: This assesses the logical flow and connectivity of the text produced by the models.
Relevance: This measures how well the model’s output adheres to the context provided in the input.
Linguistic Quality: This evaluates the grammatical accuracy, punctuation, and stylistic appropriateness of the text.
5.2.3 Evaluation Results
The evaluation results are summarized in Table 3, which provides a clear comparative analysis of the performance of the two models across all assessed criteria. The scores indicate that while Ava scored slightly higher in coherence, Co-CoLA outperformed Ava in relevance and linguistic quality, suggesting its superior ability to produce contextually accurate and linguistically refined outputs.
Criteria Co-CoLA Ava
Coherence 4.2 4.3
Relevance 3.7 3.2
Linguistic Quality 4.6 4.0
Table 3: Average Scores from Linguistic Evaluation
The higher scores of Co-CoLA in relevance and linguistic quality demonstrate its effectiveness in producing not only grammatically correct but also contextually relevant outputs, which is essential for real-world applications. These results underscore the potential of Co-CoLA in enhancing the linguistic handling of Persian language tasks, setting a benchmark for future developments in language model applications.
6 Discussion
Figure 5 provides a detailed breakdown of the overall performance reported in Table 2. Each dot in the plot represents the ROUGE-L F1 score of the given model on the selected template. As clearly illustrated, other Persian instruction-tuned models fail to achieve a high ROUGE-L F1 score. One significant factor contributing to this disparity is the low precision score. The F1 score, which combines precision and recall, serves as a comprehensive metric for evaluation. Precision measures the proportion of the longest common subsequence (LCS) in the candidate text that matches the reference text, while recall measures the proportion of the LCS in the reference text that is present in the candidate text. Although these models achieve acceptable recall scores, they fall short in precision, a critical metric for categorization templates. In contrast, Aya demonstrates proficiency in handling both generation and categorization templates within the Persian context. Compared to Aya, Co-CoLA enhances the model’s ability to manage both categorization and generation tasks effectively while being less computationally expensive.
@@ 7 Conclusion
This study introduces significant advancements with FarsInstruct and Co-CoLA, addressing critical gaps in the processing and instruction-following capabilities for Persian, a low-resource language. FarsInstruct, with its diverse tasks ranging from text summarization to named entity recognition, has proven to enhance language model performance as shown through rigorous ROUGE evaluations and human assessments. This dataset not only enriches multilingual model training but also establishes a new standard for language model instruction tuning.
Further, Co-CoLA leverages the strengths of CoLA with rehearsal training to mitigate catastrophic forgetting and improve multi-task adaptation, through its iterative optimization framework. This allows for sustained model performance over diverse tasks while optimizing computational resources. Looking ahead, the focus will be on expanding the scope of these datasets to cover more tasks and modalities, thereby driving further innovations in cross-lingual language understanding and promoting AI inclusivity.
8 Limitations
This section delineates the principal limitations of our study, which, while providing substantial contributions to Persian NLP, presents challenges that could be addressed in future developments to enhance its utility and applicability in broader linguistic contexts:
Data Diversity and Representation: Although FarsInstruct broadens the corpus of Persian language resources, it does not fully capture the rich tapestry of dialects and sociolects that characterize the Persian-speaking world. Also, the collected templates are generally biased towards short responses, which might affect the overall performance of the model.
Complexity of Instructions: The dataset prompts vary in complexity but still may not sufficiently challenge or train models to handle the types of complex instructions encountered in everyday human interactions. Real-world applications often demand a high level of interpretative depth and context awareness—qualities that current models may struggle with when trained on existing datasets. Future versions of FarsInstruct could benefit from integrating prompts that require higher-order cognitive processing, such as irony, metaphor understanding, and techniques that involve prompting the model to break down complex tasks into intermediate steps, mimicking human reasoning processes (Wei et al., 2022b).
Dependency on External Datasets: The effectiveness of the FarsInstruct dataset is contingent upon the quality and variety of the external datasets. This dependency creates vulnerability, as biases or errors in source datasets may be passed to FarsInstruct.A rigorous process for source data, coupled with efforts to develop original, high-quality training materials, could diminish reliance on external datasets and enhance the overall integrity of the dataset.
Evaluation Metrics: The metrics currently used to evaluate models trained on FarsInstruct may not fully capture the nuanced and multifaceted aspects of language comprehension and generation. Furthermore, for certain tasks such as rewriting, ROUGE-L may not serve as an adequate measure of quality.
Performance Stability: While Co-CoLA has demonstrated effectiveness in terms of computational efficiency and consistent performance across all tasks it learned, mitigating catastrophic forgetting, we observe that its overall performance is heavily dependent on the model’s performance at each tuning iteration. We leave potential solutions to this problem to future work.
Appendix
  1. A.
    Training Configuration
    All implementations were carried out using PyTorch, Transformers (Wolf et al., 2020) and Accelerate (Gugger et al., 2022) library. For efficient training, we randomly selected 25 prompt templates and applied them to their corresponding datasets. Consequently, for example, a dataset with two selected templates would be upsampled to twice its original size. We then sampled a minimum of 10,000 examples from each dataset, based on the specific template and dataset length, to create the current training data. The rehearsal hyper-parameter of Co-CoLA was set to 0.01. We used Paged-AdamW as the base optimizer and trained for a total of four epochs in each tuning phase. A linear learning rate scheduler was applied, with an initial learning rate of $6\times 10^{-5}$ and a batch size of 16. For implementing LoRA, we utilized the PEFT (Mangrulkar et al., 2022) library for convenience.
  2. B.
    Dataset Illustrations
@@ 8 Limitations
  3. C.
    Prompts
    Figure 7: A prompt example shown in Promptsource environment. PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts. Prompts function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs. In PromptSource, we develop input templates, target templates, and choice templates. Inputs typically consist of questions or instructions, while the output code specifies the expected answer or result. For Categorization tasks, the choice template includes predefined options for answering questions, while Generation tasks do not require this template. In this picture, The "Metrics" box is set to measure Accuracy for Categorization tasks, and the "Prompt Language" used is Farsi (Persian). "Answer choices" are provided within the template, which comprises an instruction followed by data fields. The premise and hypothesis are selected from the "Data Schema" on the left side of the interface. The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).

### v3

@@ 4 Methodology and Experimental Setup
To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (xia2024chain) with continual learning (kirkpatrick2017overcoming). This section offers a thorough overview of the training procedure and evaluation setup.
@@ 4.1 Training Procedure
Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset. However, as highlighted in studies by (wang2023multilora; li2024mixlora), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance. To mitigate this challenge, Chain of LoRA (CoLA) (xia2024chain), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method). This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module. xia2024chain shows that this process allows the model to learn higher-rank adaptations more effectively. Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting. kalajdzievski2024scaling observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
@@
In this study, we propose Continual-Chain of LoRA (Co-CoLA), an extension of the CoLA framework that incorporates rehearsal with replay during training. More specifically, rehearsal training is an approach within the continual learning framework that involves revisiting a portion of previously learned tasks while training new tasks. The core mathematical operation in LoRA involves updating the low-rank matrices $A$ and $B$, which are applied to modify the transformer layers of the model. The update rule can be expressed as $W^{\prime}=W+BA$ where $W$ represents the transformer layer’s original weights, and $W^{\prime}$ shows the updated weights after applying the low-rank adjustments $A$ and $B$. Essentially, Co-CoLA structures the training procedure by iterating over the following three phases:
@@
Tuning: Following the standard LoRA approach, the weights of the base model remain frozen, while only the model’s LoRA parameters, represented by matrices $A$ and $B$ are fine-tuned. During this phase, a subset of previously trained data is replayed along with the new data. Formally, let $T=(T_{1},\ldots,T_{n})$ denote the sequence where each $T_{i}$ represents the training data obtained by applying the prompt template $i$ to its corresponding dataset. the training data augmented with rehearsal is defined as:
$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$
@@
where r is the rehearsal hyperparameter that controls the percentage of examples sampled from previous tasks $T_{1},\ldots,T_{i-1}$.
@@
Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in hu2021lora. These merged weights are fixed and do not receive any gradient update in subsequent steps.
@@
Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$). Following hu2021lora $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
@@ 4.2 Evaluation Setup
The performance of our model is assessed across two categories of task types: those included in the training dataset (”Held in”) and those introduced for the first time during evaluation (”Held out”). This choice allows for a more comprehensive evaluation of the model’s generalization abilities. The evaluation dataset comprises three distinct task types: Sentiment Analysis and Query Paraphrasing, classified as “Held in” tasks, and Textual Entailment, categorized as a “Held out” task. As shown in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, as well as two datasets specifically for entailment tasks.
We employ the ROUGE-L metric to evaluate the overlap of n-grams between the generated text and reference texts. Our focus was on the F1-scores of ROUGE-L, which combines precision and recall for a comprehensive assessment. As shown by wang-etal-2022-super, the rankings generated by this metric correlate strongly with accuracy for categorization templates.
@@ 5 Results
To investigate whether FarsInstruct can enhance LLM’s performance on Persian tasks, we perform instruction tuning on an open-source Persian-English LLM as the foundation model. We additionally employ mistral, aya and llama as baselines.
In our experiments, we discovered that the FarsInstruct dataset significantly boosts the instruction-following capabilities of language models tailored for the Persian language. Initial results indicate that models trained with FarsInstruct not only improved in traditional NLP tasks like text classification and summarization but also showed enhanced performance in complex tasks such as entailment and sentiment analysis. Specifically, the use of FarsInstruct resulted in an average improvement of 15% in accuracy scores across various models when compared to those trained with existing datasets. This enhancement underscores the efficacy of FarsInstruct in addressing the gaps in data diversity and complexity for low-resource languages.
Moreover, the models exhibited increased robustness in handling nuanced linguistic features of Persian, a critical aspect often overlooked in standard datasets. During the evaluation phase, models demonstrated an adeptness in interpreting intricate instructions and generating responses that align more closely with the expectations set by native Persian speakers. The success of FarsInstruct in elevating the performance of Persian language models offers promising avenues for future research in multilingual NLP, particularly for languages that are underrepresented in the digital space.
@@ 6 Conclusion
In this work we introduce farsinstruct, a comprehensive instruction dataset specifically curated for Persian language,[bridging the critical gap in current language resources for underrepresented language and advancing LLMs training in cross lingual domain] to bridge the critical gap [sth] [what gap?]. Farsinstruct is comprised of 210 templates across 11 distinct datasets. Also to ensure the richeness of data we translated 3 popular question answering datasets. Looking ahead, we are committed to the ongoing enrichment and refinement of Farsinstruct. We will incorporate a wider array of task types, instruction entries, and modalities, in line with the latest advances in cross lingual language understanding. Moreover given the limitations imposed by Lora training strategy and computational resources current language models have yet to improve on Persian context.
V2
In concluding the study on ”FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding,” it becomes evident that this groundbreaking dataset marks a pivotal advancement in the field of NLP for low-resource languages, particularly Persian. By meticulously curating over 200 templates from a variety of sources and enriching the linguistic landscape with complex Persian tasks, FarsInstruct not only bridges a critical gap in multilingual model training but also sets a new precedent for future research in language model instruction tuning. The dataset’s comprehensive coverage, reflecting a rich array of task types and linguistic nuances, promises significant improvements in model performance and instruction following capabilities. Looking ahead, the commitment to continual updates and expansion of FarsInstruct underscores the ongoing quest for achieving greater linguistic diversity and model adaptability. This endeavor not only enriches the Persian NLP toolkit but also inspires further exploration and innovation in the domain of instruction-tuned large language models, potentially catalyzing advancements in cross-lingual understanding and global inclusivity in AI technologies.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
A.1
@@ Appendix A Appendix
A.2 Prompts
Figure 6: An example shown in the PromptSource environment. PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts. Prompt templates function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs. In PromptSource, we develop input templates, target templates, and choice templates. Inputs typically consist of questions or instructions, while the output code specifies the expected answer or result. For categorization tasks, the choice template includes predefined options for answering questions, while generation tasks do not require this template. In this picture, The ”Metrics” box is set to measure Accuracy for categorization tasks, and the ”Prompt Language” used is Farsi (Persian). ”Answer choices” are provided within the template, which comprises an instruction followed by data fields. The premise and hypothesis are selected from the ”Data Schema” on the left side of the interface. The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).